OpenStudy (anonymous):

What is the range of the relation? {(1, 2), (2, 4), (3, 2), (4, 6)} A. {2, 4, 6} B. {1, 2, 3, 4} C. {2, 4} D. {1, 2, 3, 4, 6}

OpenStudy (anonymous):

the ans-A

OpenStudy (anonymous):

thats what i thought, thanks

OpenStudy (anonymous):

one more: For the function f(x) = 2 – 3x, find f(0). A. –3 B. 5 C. 2 D. –1

OpenStudy (anonymous):

i think it easy C

OpenStudy (anonymous):

but im wondering how do u get c?

OpenStudy (the_fizicx99):

@mintefikre Refrain from giving out answers.

OpenStudy (anonymous):

can anyone tell me how mintefikre got c?

OpenStudy (the_fizicx99):

f(0) means to plug in 0 for x \(\ \sf f(\color{green}{x})\) = 2 - \(\ 3\color{green}{x} \longrightarrow 2 -3(\color{green}{0}) = 2 - 0 = 2\) Anything times 0 is 0

OpenStudy (anonymous):

Oh i didnt know what f(0) meant Thanks so much!

OpenStudy (the_fizicx99):

yw